### v1

@@ 5 Examples
In this section, we consider multiple example problems to demonstrate the applicability of the FSNN to various problems, including multi-scale problems arising in hydrodynamics, climate dynamics, and plasma physics.
@@ 5.1 Visualization of the Attracting Slow Manifold Property
A significant advantage of the FSNN is that, for any choice of model parameters, the network will have an attracting slow manifold that, due to Theorem 2, will attract solutions for sufficiently small values of $\epsilon$. The slow manifold is represented by the graph, $y=0$. In Figure 4, we demonstrate this property on multiple random initializations of FSNNs for $N_{y}=N_{x}=1$. Initial conditions are drawn randomly from a standard normal distribution (left plot) and are integrated to a later time $t=4$. The trajectory points at $t=4$ (blue dots) and the resulting slow manifold (orange curves) are plotted at this final time, demonstrating convergence of solutions to the slow manifold.
@@ 5.2 Learning a Simple Attracting Manifold
$\frac{dz_{1}}{dt}$ $=-\epsilon(\sin(z_{1})+z_{2})$ (67a)
$\frac{dz_{2}}{dt}$ $=\lambda(z_{1})(z_{2}-\theta(z_{1}))$ (67b)
@@
$\lambda(x)=-1-\frac{1}{10}\cos(2x),\qquad\theta(x)=2\tanh(x).$ (68)
@@
A FSNN is trained on this dataset using the Adam optimizer from Optax. The INN consists of one outer layer, where the inner coupling flow layers consists of feedforward networks with one layer and a hidden dimension of 10. East of the feedforward neural networks on the RHS of the neural ODE system consists of 10 layers and the bilinear form has a rank of 2.
Figure 5 shows a comparison of the trained model and the reference solution in phase space for various choices of $\epsilon$. The model shows great agreement to the reference solution. In all of the plots, it is evident that the trajectories eventually get trapped in a slow manifold. In the rightmost plot, a comparison of the learned eigenvalues with that of the ground truth, $\lambda(z_{1})$, is shown. The results demonstrate that the FSNN is able to successfully recover details of the original operator.
Figure 5: Results of learning the dynamics of the simple attracting manifold example described by (5.2). Left three plots: phase portraits comparing the ground truth (g.t.) reference solution (bold curves) and that of the trained model (dots) for $\epsilon=0,0.001,0.01$. Rightmost plot: comparison of the eigenvalues learned by the FSNN with the ground truth eigenvalues.
5.3 Grad’s Moment System
@@
$\partial_{t}\rho$ $=-\epsilon\,\nabla\cdot\mathbf{u},$ (69a)
$\partial_{t}\mathbf{u}$ $=-\epsilon\left(\nabla\rho+\nabla T+\nabla\cdot\sigma\right),$ (69b)
$\partial_{t}T$ $=-\frac{2\epsilon}{3}\left(\nabla\cdot\mathbf{u}+\nabla\cdot\mathbf{q}\right),$ (69c)
$\partial_{t}\mathbf{\sigma}$ $=-\mathbf{\sigma}-\epsilon\left(2\overline{\nabla\mathbf{u}}+\frac{4}{5}\overline{\nabla\mathbf{q}}\right),$ (69d)
$\partial_{t}\mathbf{q}$ $=-\frac{2}{3}\mathbf{q}-\epsilon\left(\frac{5}{2}\nabla T+\mathbf{\nabla}\cdot\mathbf{\sigma}\right),$ (69e)
@@
$\overline{\mathbf{a}}$ $=\frac{1}{2}(\mathbf{a}+\mathbf{a}^{T})-\frac{1}{3}\mathbf{I}\,\text{Tr}(\mathbf{a}).$ (70)
@@
$\rho(\mathbf{x},t)=\sum_{k=-\infty}^{\infty}\rho_{k}(t)e^{ikx}.$ (71)
@@
$\frac{d}{dt}\left[\begin{array}[]{c}\rho_{k}\\ u_{k}\\ T_{k}\\ \sigma_{k}\\ q_{k}\end{array}\right]=\left[\begin{array}[]{ccccc}0&-ik\epsilon&0&0&0\\ -ik\epsilon&0&-ik\epsilon&-ik\epsilon&0\\ 0&-\frac{2}{3}ik\epsilon&0&0&-\frac{2}{3}ik\epsilon\\ 0&-\frac{4}{3}ik\epsilon&0&-1&-\frac{8}{15}ik\epsilon\\ 0&0&-\frac{5}{2}ik\epsilon&-ik\epsilon&-\frac{2}{3}\end{array}\right]\left[\begin{array}[]{c}\rho_{k}\\ u_{k}\\ T_{k}\\ \sigma_{k}\\ q_{k}\end{array}\right],\quad k=0,\pm 1,\dots$ (87)
For this example, we truncate the Fourier series to consider only the $k=0,\pm 1,\pm 2$ models. Initial conditions are randomly generated for each Fourier coefficient uniformly in the complex region $[-1,1]\times[-i,i]$. The solution is computed for 10 time steps both using $\Delta t_{\rm fast}=\frac{1}{4}$ and $\Delta t_{\rm slow}=\frac{1}{4\epsilon}$. The slow-scale trajectories are offset $\Delta t_{\rm slow}$ from $t=0$ to ensure the solution is sufficiently close to the slow manifold. A FSNN is trained on this dataset using the Adam optimizer from Optax. The INN consists of one outer layer, where the inner coupling flow layers consists of feedforward networks with one layer and a hidden dimension of 40. The feedforward neural networks on the RHS of the neural ODE system each consist of 50 layers and the bilinear form has a rank of 1. Considering each coefficient and its real and imaginary parts, the dimension of the learned model is $\mathbb{R}^{30}$, with $N_{x}=18$ and $N_{y}=12$.
Table 1: Comparison of ground truth eigenvalues of the Jacobian of the fast equations with that of the trained model for the Grad Moment System problem.
@@
Figure 6 shows comparisons of the trained model and reference solutions on the fast time scale and slow time scale. The results demonstrate that the model is accurate on both time scales. Table 1 shows a comparison of the eigenvalues of the trained model with the analytical eigenvalues. Great agreement is observed, which implies that the model was successful in mapping the slow manifold approximately to the true slow manifold.
It is critical to stress that the slow dynamics comparison in Figure 6 is done through integration on the slow manifold using (3.4). Specifically, the initial condition state vector $z_{0}$ is first transformed into the fast-slow coordinates $(x_{0},y_{0})$ and projected onto the learned slow manifold $(x_{0},0)$ before integrating the slow variable $x$ on the slow time scale only and then transforming back into original coordinates using $h^{-1}$. This approach is significantly more efficient than integrating the full system. The excellent accuracy of the predicted slow variable solution indicates a data-driven closure is learned through this procedure. The closure of (5.3) formally exists when $\epsilon$ approaches 0. It is readily seen that the reduced system is a linearized Euler system.
@@ 5.4 Multi-Scale Lorenz96 Equations
In this section we consider the two-scale Lorenz96 equations, which was a model proposed in [47] as a test bed model for multi-scale behavior in atmospheric dynamics. The original system is formulated as
$\frac{d}{d\tau}X_{k}$ $=-X_{k-1}(X_{k-2}-X_{k+1})-X_{k}+F-(hc/b)\sum_{j=1}^{J}Y_{j,k},$ (88a)
$\frac{d}{d\tau}Y_{j,k}$ $=-cbY_{j+1,k}(Y_{j+2,k}-Y_{j-1,k})-cY_{j,k}+(hc/b)X_{k},$ (88b)
@@
$\frac{d}{dt}x_{k}$ $=-\epsilon x_{k-1}(x_{k-2}-x_{k+1})-\epsilon x_{k}-\epsilon^{2}\sum_{j=1}^{J}y_{j,k},$ (89a)
$\frac{d}{dt}y_{j,k}$ $=-y_{j+1,k}(y_{j+2,k}-y_{j-1,k})-y_{j,k}+x_{k}.$ (89b)
@@
$(D_{y}f)_{jk}=\frac{\partial f_{j}}{\partial y_{k}}=-\delta_{jk}-\delta_{{j+1},k}(y_{j+2}-y_{j-1})-y_{j+1}(\delta_{j+2,k}-\delta_{j-1,k}),$ (91)
@@
$(D_{y}f)_{jk}|_{y=\xi}=-\delta_{jk}-\chi(\delta_{j+2,k}-\delta_{j-1,k})$ (92)
@@
$\lambda_{j}=-1-\chi(e^{ij4\pi/J}-e^{-ij2\pi/J}),\qquad j=0,1,\dots,J-1.$ (93)
@@
The equations defined by (5.4) are energy-stable for $\epsilon>0$.
@@
$x_{k}x_{k}^{\prime}$ $=-\epsilon x_{k}x_{k-1}(x_{k-2}-x_{k+1})-\epsilon x_{k}^{2}-\epsilon^{2}\sum_{j=1}^{J}x_{k}y_{j,k}$ (95a)
$y_{j,k}y_{j,k}^{\prime}$ $=-y_{j,k}y_{j+1,k}(y_{j+2,k}-y_{j-1,k})-y_{j,k}^{2}+x_{k}y_{j,k}.$ (95b)
@@
$\frac{1}{2}\sum_{k=1}^{K}(x_{k}^{2})^{\prime}+\frac{1}{2}\epsilon^{2}\sum_{k=1}^{K}\sum_{j=1}^{J}(y_{j,k}^{2})^{\prime}=-\epsilon\sum_{k=1}^{K}x_{k}^{2}-\epsilon^{2}\sum_{k=1}^{K}\sum_{j=1}^{J}y_{j,k}^{2}<0$ (96)
We train a FSNN to learn the multi-scale Lorenz96 equations from data. The trajectory dataset, $\mathcal{D}$, consists of $10^{4}$ fast-scale trajectories, corresponding to initial conditions for $x_{k}$ randomly generated uniformly in $[-\frac{1}{2},1]$ and $y_{j,k}$ randomly generated from a normal distribution, $\epsilon$ randomly generated log-uniformly in $[10^{-5},10^{-2}]$, and integrated forward 10 timesteps using an adaptive RK45 solver with error tolerance $10^{-8}$. The manifold dataset, $\mathcal{M}$ consists of $10^{4}$ examples of data pairs $((x_{k},y_{j,k}),\epsilon=0)$ on the graph $y_{j,k}=x_{k}$, where $x_{k}$ is randomly generated uniformly in $[-\frac{1}{2},1]$.
A FSNN is trained on this dataset using the Adam optimizer from Optax. The INN consists of one outer layer, where the inner coupling flow layers consists of feedforward networks with one layer and a hidden dimension of 40. The feedforward neural networks on the RHS of the neural ODE system each consist of 40 layers and the bilinear form has a rank of 4.
Figure 7: The top five plots are trajectory plots comparing the reference solution computed using RK45 (bold curves) and that of the trained model (dotted markers) for an initial condition with $\epsilon=1.1\cdot 10^{-3}$ for the multi-scale Lorenz96 example described by the dynamics in (5.4). The bottom five plots show the corresponding errors as a function of time. The plots in the left column show the solution and error on the slow time scale ($t\in[0,3000]$) while the plots in the right column zoom in on the fast time scale ($t\in[0,3]$).
Figure 7 shows a comparison of the trained model and the reference solution for an example initial condition with $\epsilon=1.1\cdot 10^{-3}$. The model shows quantitative agreement to the reference solution. Figure 8 shows a comparison of the eigenvalues of the trained model with the analytical eigenvalues. Agreement is observed, implying that the model was successful in mapping the slow manifold approximately to the true slow manifold. In this example, the time interval of all the training data is much shorter than $t=3$. Nevertheless, the prediction can be done up to $T=3000$ or longer with excellent accuracy.
Figure 8: Comparison of the true eigenvalues (bold lines) with those learned from the trained model (dots) for the multi-scale Lorenz96 example described by the dynamics in (5.4)

### v2

@@ 5 Examples
In this section, we consider multiple example problems to demonstrate the applicability of the FSNN to various problems, including multi-scale problems arising in hydrodynamics and plasma physics. In training, we explored various combinations of meta parameters for each problem. We pursued a policy to train several models in parallel and choose the model with the best performance. The meta parameters for the best models are reported in each respective section. We remark that though increasing the number of network parameters increases the expressiveness of the model, it comes at a higher computational cost and increases the risk of overfitting.
@@ 5.1 Visualization of the Attracting Slow Manifold Property
A significant advantage of the FSNN is that, for any choice of model parameters, the network will have an attracting slow manifold that, due to Theorem 8, will attract solutions for sufficiently small values of $\epsilon$. The slow manifold is represented by the graph, $y=0$. In Figure 4, we demonstrate this property on multiple random initializations of FSNNs for $N_{y}=N_{x}=1$. In this figure, network a, b, and c are representing different FSNNs with different weights and therefore different slow manifolds. Multiple initial conditions are drawn randomly from a standard normal distribution (left plot) and are integrated using each FSNN to a later time $t=4$. The trajectory points at $t=4$ (blue dots) and the resulting slow manifold (orange curves) are plotted at this final time, demonstrating convergence of solutions to the slow manifolds inherent to each network.
@@ 5.2 Learning a Simple Attracting Manifold
(66a) $\frac{dz_{1}}{dt}$ $=-\epsilon(\sin(z_{1})+z_{2})$
(66b) $\frac{dz_{2}}{dt}$ $=\lambda(z_{1})(z_{2}-\theta(z_{1}))$
@@
(67) $\lambda(x)=-1-\frac{1}{10}\cos(2x),\qquad\theta(x)=2\tanh(x).$
@@
A FSNN is trained on this dataset using the Adam optimizer from Optax. The INN consists of one outer layer, where the inner coupling flow layers consists of feedforward networks with one layer and a hidden dimension of 10. Each of the feedforward neural networks on the RHS of the neural ODE system consists of 10 layers and the bilinear map has a rank of 2.
Figure 5 shows a comparison of the trained model and the reference solution in phase space for various choices of $\epsilon$. As can be seen by the corresponding error plots, the model shows reasonable agreement to the reference solution. In all of the phase plots, it is evident that the trajectories eventually get trapped in a slow manifold. In the rightmost plot on the top row, a comparison of the learned eigenvalues with that of the ground truth, $\lambda(z_{1})$, is shown. Additionally in the rightmost plot on the bottom row, the error between the learned slow manifold and true slow manifold is shown. The results demonstrate that the FSNN is able to successfully recover details of the original operator.
Figure 5: Results of learning the dynamics of the simple attracting manifold example described by (66). Left three plots on top row: phase portraits comparing the ground truth (g.t.) reference solution (bold curves) and that of the trained model (dots) for $\epsilon=0,0.001,0.01$. Left three plots on bottom row: corresponding errors in time for each trajectory in the above phase portraits. Rightmost plot on top row: comparison of the eigenvalue of the negative Schur form as a function of $z_{1}$ learned by the FSNN with the ground truth eigenvalue given by $\lambda$ in (67). Rightmost plot on bottom row: error between the learned slow manifold and the ground truth slow manifold given by $\theta$ in (67).
5.3 Linear Grad Moment System
@@
(68a) $\partial_{t}\rho$ $=-\epsilon\,\nabla\cdot\mathbf{u},$
(68b) $\partial_{t}\mathbf{u}$ $=-\epsilon\left(\nabla\rho+\nabla T+\nabla\cdot\sigma\right),$
(68c) $\partial_{t}T$ $=-\frac{2\epsilon}{3}\left(\nabla\cdot\mathbf{u}+\nabla\cdot\mathbf{q}\right),$
(68d) $\partial_{t}\mathbf{\sigma}$ $=-\mathbf{\sigma}-\epsilon\left(2\overline{\nabla\mathbf{u}}+\frac{4}{5}\overline{\nabla\mathbf{q}}\right),$
(68e) $\partial_{t}\mathbf{q}$ $=-\frac{2}{3}\mathbf{q}-\epsilon\left(\frac{5}{2}\nabla T+\mathbf{\nabla}\cdot\mathbf{\sigma}\right),$
@@
(69) $\overline{\mathbf{a}}$ $=\frac{1}{2}(\mathbf{a}+\mathbf{a}^{T})-\frac{1}{3}\mathbf{I}\,\text{Tr}(\mathbf{a}).$
@@
(70) $\rho(\mathbf{x},t)=\sum_{k=-\infty}^{\infty}\rho_{k}(t)e^{ikx}.$
@@
(86) $\frac{d}{dt}\left[\begin{array}[]{c}\rho_{k}\\ u_{k}\\ T_{k}\\ \sigma_{k}\\ q_{k}\end{array}\right]=\left[\begin{array}[]{ccccc}0&-ik\epsilon&0&0&0\\ -ik\epsilon&0&-ik\epsilon&-ik\epsilon&0\\ 0&-\frac{2}{3}ik\epsilon&0&0&-\frac{2}{3}ik\epsilon\\ 0&-\frac{4}{3}ik\epsilon&0&-1&-\frac{8}{15}ik\epsilon\\ 0&0&-\frac{5}{2}ik\epsilon&-ik\epsilon&-\frac{2}{3}\end{array}\right]\left[\begin{array}[]{c}\rho_{k}\\ u_{k}\\ T_{k}\\ \sigma_{k}\\ q_{k}\end{array}\right],\quad k=0,\pm 1,\dots$
For this example, we truncate the Fourier series to consider only the $k=0,\pm 1$ models. Initial conditions are randomly generated for each Fourier coefficient uniformly in the complex region $[-1,1]\times[-i,i]$. The solution is computed for 50 time steps both using $\Delta t_{\rm fast}=\frac{1}{4}$ and $\Delta t_{\rm slow}=\frac{1}{4\epsilon}$. The model is trained using only the information from the first time step and tested using the remaining of the trajectory. The slow-scale trajectories are offset $\Delta t_{\rm slow}$ from $t=0$ to ensure the solution is sufficiently close to the slow manifold. A FSNN is trained on this dataset using the Adam optimizer from Optax. The INN consists of one outer layer, where the inner coupling flow layers consists of feedforward networks with one layer and a hidden dimension of 40. The feedforward neural networks on the RHS of the neural ODE system each consist of 50 layers and the bilinear map has a rank of 1. Considering each coefficient and its real and imaginary parts, the dimension of the learned model is $\mathbb{R}^{30}$, with $N_{x}=18$ and $N_{y}=12$.
Table 1: Comparison of ground truth eigenvalues of the Jacobian of the fast equations with that of the negative Schur form of the trained model for the Grad Moment System problem.
@@
Figure 6 shows comparisons of the trained model and reference solutions on the fast time scale and slow time scale. The results demonstrate that reasonable accuracy can be achieved on both time scales. We observe that model for the slow dynamics is slightly off in phase which causes error to grow in time.
Table 1 shows a comparison of the eigenvalues of the trained model with the analytical eigenvalues. Quantitative agreement is observed, which implies that the model was successful in mapping the slow manifold approximately to the true slow manifold.
It is critical to stress that the slow dynamics comparison in Figure 6 is done through integration on the slow manifold using (61). Specifically, the initial condition state vector $z_{0}$ is first transformed into the fast-slow coordinates $(x_{0},y_{0})$ and projected onto the learned slow manifold $(x_{0},0)$ before integrating the slow variable $x$ on the slow time scale only and then transforming back into original coordinates using $h^{-1}$. This approach is significantly more efficient than integrating the full system using a standard numerical integrator. The excellent accuracy of the predicted slow variable solution indicates a data-driven closure is learned through this procedure. The closure of (68) formally exists when $\epsilon$ approaches 0. It is readily seen that the reduced system is a linearized Euler system. When $\epsilon=0$, the critical manifold is simply the center subspace. When weak nonlinearity is added, there are center manifolds tangent to the center subspace. In general these center manifolds are not unique, but there is a unique smoothest center manifold known as the spectral submanifold, see [64, 65]. In future work, it would be interesting to explore whether our methods can be adapted to ensure the learned slow manifold agrees with the spectral submanifold.
@@ 5.4 Multi-Scale Lorenz96 Equations
In this section we consider the two-scale Lorenz96 equations, which was a model proposed in [66] as a test bed model for multi-scale behavior in atmospheric dynamics. Multiscale techniques for this system have been explored in many works, such as in [67]. The original system is formulated as
(87a) $\frac{d}{d\tau}X_{k}$ $=-X_{k-1}(X_{k-2}-X_{k+1})-X_{k}+F-(hc/b)\sum_{j=1}^{J}Y_{j,k},$
(87b) $\frac{d}{d\tau}Y_{j,k}$ $=-cbY_{j+1,k}(Y_{j+2,k}-Y_{j-1,k})-cY_{j,k}+(hc/b)X_{k},$
@@
(88a) $\frac{d}{dt}x_{k}$ $=-\epsilon x_{k-1}(x_{k-2}-x_{k+1})-\epsilon x_{k}-\epsilon^{2}\sum_{j=1}^{J}y_{j,k},$
(88b) $\frac{d}{dt}y_{j,k}$ $=-y_{j+1,k}(y_{j+2,k}-y_{j-1,k})-y_{j,k}+x_{k}.$
Since the transformation is singular at $\epsilon=0$, we study the equations in the region $0<\epsilon\ll 1$.
@@
(90) $(D_{y}f)_{jk}=\frac{\partial f_{j}}{\partial y_{k}}=-\delta_{jk}-\delta_{{j+1},k}(y_{j+2}-y_{j-1})-y_{j+1}(\delta_{j+2,k}-\delta_{j-1,k}),$
@@
(91) $(D_{y}f)_{jk}|_{y=\xi}=-\delta_{jk}-\chi(\delta_{j+2,k}-\delta_{j-1,k})$
@@
(92) $\lambda_{j}=-1-\chi(e^{ij4\pi/J}-e^{-ij2\pi/J}),\qquad j=0,1,\dots,J-1.$
@@
The equations defined by (88) are energy-stable for $\epsilon>0$.
@@
(94a) $x_{k}x_{k}^{\prime}$ $=-\epsilon x_{k}x_{k-1}(x_{k-2}-x_{k+1})-\epsilon x_{k}^{2}-\epsilon^{2}\sum_{j=1}^{J}x_{k}y_{j,k}$
(94b) $y_{j,k}y_{j,k}^{\prime}$ $=-y_{j,k}y_{j+1,k}(y_{j+2,k}-y_{j-1,k})-y_{j,k}^{2}+x_{k}y_{j,k}.$
@@
(95) $\frac{1}{2}\sum_{k=1}^{K}(x_{k}^{2})^{\prime}+\frac{1}{2}\epsilon^{2}\sum_{k=1}^{K}\sum_{j=1}^{J}(y_{j,k}^{2})^{\prime}=-\epsilon\sum_{k=1}^{K}x_{k}^{2}-\epsilon^{2}\sum_{k=1}^{K}\sum_{j=1}^{J}y_{j,k}^{2}<0$
We train a FSNN to learn the multi-scale Lorenz96 equations from data choosing $J=K=4$. The trajectory dataset, $\mathcal{D}$, consists of $10^{4}$ fast-scale trajectories, corresponding to initial conditions for $x_{k}$ randomly generated uniformly in $[-\frac{1}{2},1]$ and $y_{j,k}$ randomly generated from a normal distribution, $\epsilon$ randomly generated log-uniformly in $[10^{-5},10^{-2}]$, and integrated forward 10 timesteps using an adaptive RK45 solver with error tolerance $10^{-8}$. The manifold dataset, $\mathcal{M}$ consists of $10^{4}$ examples of data pairs $((x_{k},y_{j,k}),\epsilon=0)$ on the graph $y_{j,k}=x_{k}$, where $x_{k}$ is randomly generated uniformly in $[-\frac{1}{2},1]$.
A FSNN is trained on this dataset using the Adam optimizer from Optax. The INN consists of one outer layer, where the inner coupling flow layers consists of feedforward networks with one layer and a hidden dimension of 40. The feedforward neural networks on the RHS of the neural ODE system each consist of 40 layers and the bilinear map has a rank of 4.
Figure 7: The top five plots are trajectory plots comparing the reference solution computed using RK45 (bold curves) and that of the trained model (dotted markers) for an initial condition with $\epsilon=1.1\cdot 10^{-3}$ for the multi-scale Lorenz96 example described by the dynamics in (88). The bottom five plots show the corresponding errors as a function of time. The plots in the left column show the solution and error on the slow time scale ($t\in[0,3000]$) while the plots in the right column zoom in on the fast time scale ($t\in[0,3]$).
Figure 7 shows a comparison of the trained model and the reference solution for an example initial condition with $\epsilon=1.1\cdot 10^{-3}$. The model shows quantitative agreement to the reference solution. Although the globally attracting solution is the zero solution, the error remains away from 0. This can be explained by approximation error of the model and can be improved through including more data in the training set. Figure 8 shows a comparison of the eigenvalues of the trained model with the analytical eigenvalues. Agreement is observed, implying that the model was successful in mapping the slow manifold approximately to the true slow manifold. In this example, the time interval of all the training data is much shorter than $t=3$. Nevertheless, the prediction can be done up to $T=3000$ or longer with excellent accuracy.
Figure 8: Comparison of the ground truth eigenvalues of the Jacobian of the fast dynamics (bold lines) with that of the negative Schur form learned from the trained model (dots) for the multi-scale Lorenz96 example described by the dynamics in (88)